1. H > (M > E)
2. ~[(M > H) > E]
3. (I + O) > [M + (X = P)]
4. [(K v B) > ~M] > (~I v ~Z)
5. [I > ~(Z v O)] > (A > ~K)
6. E v (K v A) / A = ~K

Answer by CPhill(1959)

To analyze the given logical expressions, we can use a truth table or apply logical rules. Let's use the latter approach:
Given:
H > (M > E)
~[(M > H) > E]
(I + O) > [M + (X = P)]
[(K v B) > ~M] > (~I v ~Z)
[I > ~(Z v O)] > (A > ~K)
E v (K v A) / A = ~K
Analysis:
H > (M > E): This means "If H is true, then if M is true, E must also be true."
~[(M > H) > E]: This is equivalent to "(M > H) and ~E." In other words, "M implies H, but E is false."
(I + O) > [M + (X = P)]: This means "If both I and O are true, then either M is true or X equals P."
[(K v B) > ~M] > (~I v ~Z): This means "If either K or B implies not-M, then either not-I or not-Z must be true."
[I > ~(Z v O)] > (A > ~K): This means "If I implies neither Z nor O, then A implies not-K."
E v (K v A) / A = ~K: This is a conditional statement with a conclusion. It means "If either E is true, or K or A is true, then A implies not-K."
Conclusion:
Without specific truth values for the variables, we cannot definitively determine the overall truth value of the entire expression. However, we can analyze the implications of each statement and how they interact with each other.
To further analyze and draw specific conclusions, we would need more information, such as specific truth assignments for the variables or additional constraints on the system.
Note: The symbol "+" is often used to denote logical OR, and "v" is often used to denote logical OR as well. The symbol "=" is often used to denote logical equivalence.